Calculate the amount (in moles) of (a) Cu2+ ions in 3.00 g of CuBr2
Answer #1

Molar mass of CuBr2 = 63.546+ 2×79.904=223.35 g/mol.

This means 223.35 g CuBr2 = 1 mol CuBr2 ,

then 1 g of CuBr2 = (1/223.35) mol CuBr2

So 3.00 g CuBr2 = (1/223.35)×3.00 mol CuBr2 = 0.0134 mol CuBr2

In 1 CuBr2 there is one Cu2+ ion.

Thus moles of Cu2+ ions in 3.00 g CuBr2 is 0.0134 mol.

Comment in case of any doubt.